2/7 of a week = 2
how do you get it

Found 2 solutions by ewatrrr, solver91311:
Answer by ewatrrr(24785)

 
Hi
Note: Measure of 'anything' needs to be accompanied by the unit of measurement.
2/7 of a week = 2 days ||week = 7 days, (2/7)7days = 2 days

Answer by solver91311(24713)



The expression is not necessarily true. 2/7 of a week = 2 days IS true. But 2/7 of a week = 2 (any time period other than a day) is a false statement.

A fraction can be considered in this way: The denominator is the number of pieces into which the whole is divided, and the numerator is the number of those pieces you have. So if the time period above is in fact days, then there are 7 days in a week and that is the source of the denominator. So if you divide a week into 7 equal pieces, each one is a day, and 2 of those pieces make 2 days.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism